The Omega Company manufactures customized motors on a job-order basis.

Job 492 is an order for 100 units. It requires the following: ​

direct materials $3,000

direct labor ($10/hr) 1,000

overhead (150% DL$) 1,500 ​

After inspection, 2 units required rework which required 8 additional direct labor hours and $60 of materials.

If the rework is considered abnormal spoilage, what is the cost of job 492?
Answer #1

Total cost before spoilage

= Direct Materials + Direct Labor + overhead

= 3000+1000+1500

= 5500

Cost for Rework = Direct Materials 60 + Direct Labor (8*10) + Overhead (8*10*150%)

= 60 + 80 + 120 = 260

Total Cost of job = 5500 + 260

= $5760